# Measuring the Evolution of the NuSTAR Detector Gains

### v2

@@ 1 Introduction
The conversion of the pulse heights (PHAs) collected by the NuSTAR detectors is a multi-stage process that relies on several underlying CALDB files (for a full description see the NuSTAR data analysis software users guide). Several corrections were made to the “charge-loss correction” (CLC) files after launch to account for differences in the gain for the FPMA detectors compared to pre-launch estimates using the in-flight ${}^{155}$u calibration source. The primary gamma-ray decay lines are at high energy (86.54 and 105.4 keV), while there are several blended X-ray lines at low energy (including strong lines at 6.06 and 6.71 keV). In Madsen et al. (2015) we used two epochs of the calibration source just after launch in 2012 and in January 2015 to determine the time-dependent change in the energy scale using the standard xspec  formalism, where the transfer function to the correct PI values is:
@@ 2 Monitoring long-term gain variations
Data for each detector are integrated over six-month intervals, at which point we clean the data by requiring that the telescope bore-sight is more than ${}^{\circ}$ below the horizon to ensure that strong sources are fully blocked by the Earth. Standard OCCULTED data files include times when the bore-sight is ${}^{\circ}$ above the horizon, which is a conservative limit to avoid any attenuation in the Earth’s atmosphere. In addition, we require that the telescope is in Earth shadow to avoid any impact for solar activity.
@@
The line components are fit using several gauss components, some of which phenomenologically fit to the data (for $E>150$ keV) and some of which are known to originate from the radioactive ${}^{155}$u source. The ones primarily used for our analysis have line energies of 105.4 keV and 144.6 keV. There are also strong lines at lower energies, but these are contaminated by internal activation lines (which are also time-variable) in the CZT detectors and are not usable for this analysis.
@@
Table 1: The epochs, CALDB version numbers, and the applied changes over time at each epoch starting with the v007 CALDB file. **held fixed
@@ 3 Offset Monitoring
Figure 6: 3–20 keV images from FPMA of Kepler for the observations listed in Table 2. Observations used in the model fits are shown in green. Epoch 22 has not yet been included in this figure.
@@
14 40001020002 2014-10-11 246 ks ${}^{\circ}$ No
17a 90201021002 2017-02-07 126 ks ${}^{\circ}$ No
17b 90201021004 2017-04-22 51 ks ${}^{\circ}$ Yes
17c 90201021006 2017-10-08 54 ks ${}^{\circ}$ No
18a 90201021008 2018-06-04 42 ks ${}^{\circ}$ Yes
18b 90201021010 2018-06-17 47 ks ${}^{\circ}$ No
19 10501005002 2019-03-17 100 ks ${}^{\circ}$ Yes
20 10601005002 2020-03-18 100 ks ${}^{\circ}$ Yes
21a 10701407002 2021-05-04 18 ks ${}^{\circ}$ No
21b 10701407004 2021-05-10 50 ks ${}^{\circ}$ Yes
22 10801407002 2022-04-27 100 ks ${}^{\circ}$ Yes
23 10901407002 2023-05-04 112 ks ${}^{\circ}$ Yes
@@ Appendix A History of NuSTAR Gain-Related CALDB Files:
    Released in March 2015 and the first data set based on the 2015 ${}^{155}$u source deployment. The CLC file was update based on revised analysis of 2012 data and time-dependent Slope and Offset values were stored in the gain file.
@@ Appendix B Internal list of gain files
    FPMA DET2 gain offset at launch ($\rightarrow$ released as version 008)
@@
    2021 gain updates ($\rightarrow$ released as version 009)
@@
    NOTE: As of this release, the versioning for FPMA and FPMB will be different. The latest version of FPMB will remain at v009 while the latest on FPMA will be v010
    Contains updates to the FPMA DET2 gain described above ($\rightarrow$ released as version 010 for FPMA only)
@@
    Contains updates to the FPMA DET0 gain described above ($\rightarrow$ released as version 011 for FPMA only)

### v3

@@ 1 Introduction
The conversion of the pulse heights (PHAs) collected by the NuSTAR detectors is a multi-stage process that relies on several underlying CALDB files (for a full description see the NuSTAR data analysis software users guide). Several corrections were made to the “charge-loss correction” (CLC) files after launch to account for differences in the gain for the FPMA detectors compared to pre-launch estimates using the in-flight 155Eu calibration source. The primary gamma-ray decay lines are at high energy (86.54 and 105.4 keV), while there are several blended X-ray lines at low energy (including strong lines at 6.06 and 6.71 keV). In Madsen et al. (2015) we used two epochs of the calibration source just after launch in 2012 and in January 2015 to determine the time-dependent change in the energy scale using the standard xspec  formalism, where the transfer function to the correct PI values is:
@@ 2 Monitoring long-term gain variations
Data for each detector are integrated over six-month intervals, at which point we clean the data by requiring that the telescope bore-sight is more than 3 below the horizon to ensure that strong sources are fully blocked by the Earth. Standard OCCULTED data files include times when the bore-sight is 3 above the horizon, which is a conservative limit to avoid any attenuation in the Earth’s atmosphere. In addition, we require that the telescope is in Earth shadow to avoid any impact for solar activity.
@@
The line components are fit using several gauss components, some of which phenomenologically fit to the data (for $E>150$ keV) and some of which are known to originate from the radioactive 155Eu source. The ones primarily used for our analysis have line energies of 105.4 keV and 144.6 keV. There are also strong lines at lower energies, but these are contaminated by internal activation lines (which are also time-variable) in the CZT detectors and are not usable for this analysis.
@@
Table 1: The epochs, CALDB version numbers, and the applied changes over time at each epoch starting with the v007 CALDB file.
**held fixed
@@ 2.4 Early 2024 Update
2.5 Late 2024 Update
We tracked the apparent energy of the line centroids through $\approx$ November of 2024 and found that none of the detectors required an additional data point.
@@ 3 Offset Monitoring
Figure 6: 3–20 keV images from FPMA of Kepler for the observations listed in Table 2. Observations used in the model fits are shown in green. Epochs 22 and later were obtained at approximately the same roll angle as Epoch 21b and are omitted from this figure.
@@
14 40001020002 2014-10-11 246 ks 336 No
17a 90201021002 2017-02-07 126 ks 158 No
17b 90201021004 2017-04-22 51 ks 156 Yes
17c 90201021006 2017-10-08 54 ks 336 No
18a 90201021008 2018-06-04 42 ks 147 Yes
18b 90201021010 2018-06-17 47 ks 6.2 No
19 10501005002 2019-03-17 100 ks 158 Yes
20 10601005002 2020-03-18 100 ks 157 Yes
21a 10701407002 2021-05-04 18 ks 155 No
21b 10701407004 2021-05-10 50 ks 155 Yes
22 10801407002 2022-04-27 100 ks 154 Yes
23 10901407002 2023-05-04 112 ks 155 Yes
24 11001407002 2024-04-17 95 ks 155 Yes
@@ Appendix A History of NuSTAR Gain-Related CALDB Files:
    Released in March 2015 and the first data set based on the 2015 155Eu source deployment. The CLC file was update based on revised analysis of 2012 data and time-dependent Slope and Offset values were stored in the gain file.
@@ Appendix B Internal list of gain files
    FPMA DET2 gain offset at launch ($\rightarrow$ released as version 008 in CALDB 20191213)
@@
    2021 gain updates ($\rightarrow$ released as version 009 in CALDB 20211020)
@@
    NOTE: As of this release, the versioning for FPMA and FPMB will be different. Contains updates to the FPMA DET2 gain described above ($\rightarrow$ released as version 010 for FPMA only) in CALDB 20220510.
@@
    Contains updates to the FPMA DET0 gain described above ($\rightarrow$ released as version 011 for FPMA only) released in CALDB 20240229.